### v1

# Mid-Infrared Dust Evolution and Late-time Circumstellar Medium Interaction in SN 2017eaw

Abstract
We present JWST/MIRI and complementary ground-based near-infrared observations of the Type II SN 2017eaw taken 6 years post-explosion. SN 2017eaw is still detected out to 25 m and there is minimal evolution in the mid-infrared spectral energy distribution (SED) between the newly acquired JWST/MIRI observations and those taken a year earlier. Modeling of the mid-infrared SED reveals a cool 160 K dust component of and a hot 1700 K component of both composed of silicate dust. Notably there is no evidence of temperature or mass evolution in the cool dust component in the year between JWST observations. We also present new and archival HST and ground-based ultraviolet (UV) and optical observations which reveal reduced but continued circumstellar medium (CSM)-ejecta interaction at 2000 days post-explosion. The UV and mid-infrared emission show similar decline rates, suggesting both probe the interface between the ejecta and CSM. Given this, the continued existence of boxy H emission in the nebular spectra, the low inferred optical depth of the dust, and the lack of temperature and mass evolution, we suggest that the cool dust component in SN 2017eaw may be primarily due to pre-existing dust rather than newly-formed dust in the ejecta or cold dense shell.
1 Introduction
Observations of high redshift galaxies have revealed significant amounts of dust in the early universe (D. P. Marrone et al., 2018; T. Hashimoto et al., 2019; J. Witstok et al., 2023; V. Markov et al., 2024; A. Nanni et al., 2025). The majority of this dust is likely associated with core-collapse supernovae (CCSNe; H. L. Morgan & M. G. Edmunds, 2003; R. Maiolino et al., 2004; C. Gall et al., 2011; R. Schneider & R. Maiolino, 2024). Models of high redshift supernovae (SNe) and star formation rates indicate that SNe would need to produce between M⊙ of dust per SN (P. Todini & A. Ferrara, 2001; A. Sarangi et al., 2018; R. Schneider & R. Maiolino, 2024) and observations of nearby SN remnants have revealed dust masses within this range (L. Dunne et al., 2003; H. L. Morgan et al., 2003; J. Rho et al., 2008; I. De Looze et al., 2017; J. Chastenet et al., 2022; F. D. Priestley et al., 2022). However, the vast majority of near- and mid-infrared studies of nearby SNe undertaken prior to JWST have revealed significantly lower dust masses (10-2 ; C. Gall et al., 2011; T. Szalai & J. Vinkó, 2013; T. Szalai et al., 2019a; S. Tinyanont et al., 2016). These previous studies therefore suggest that infrared (IR) observations of CCSNe in the decades after explosion may be missing a significant portion of the dust.
Dust formation in SNe likely occurs in the expanding ejecta interior to the reverse shock and/or in a cold dense shell between the forward and reverse shock created by the interaction between the forward shock and surrounding dense circumstellar material (CSM) (e.g. M. Pozzo et al., 2004; S. Mattila et al., 2008; N. Smith et al., 2008, 2009). In the decades to centuries following the SN, some of the dust in the interior ejecta will be destroyed by interaction with the reverse shock; however, some percentage is expected to survive this interaction and facilitate further dust formation. There are some indications that the majority of newly formed dust in the first years post-explosion in non-interacting type II supernovae (SNe II111We use SNe II to refer only to SNe IIP/L SNe and exclude SNe IIb and IIn. However, we do include peculiar SNe II, such as SN 1987A-like objects, in this class.) is interior and optically thick and thus only visible in the spectral line profiles and not in infrared images. Studies modeling the optical nebular spectra of SN II recover higher dust masses than indicated by infrared photometry alone (M. Niculescu-Duvaz et al., 2022; S. Zsíros et al., 2024). Further, the significant dust mass in SN 1987A is only observable in the far-IR and sub-millimeter (P. Bouchet et al., 2006; M. Matsuura et al., 2011; R. Indebetouw et al., 2014; M. Matsuura et al., 2015; P. Cigan et al., 2019). Thus even in the case where the dust is optically thin, any newly-formed dust may be too cold to detect in the near/mid-infrared.
Measuring dust formation in SNe is further complicated by the presence of pre-existing dust in the CSM. Dust directly around the progenitor is destroyed immediately following the explosion but pre-SN dust can survive at further distances. This pre-existing dust can be formed within the CSM in the stellar winds and/or binary interactions of massive stars. This dust is warmed by the SN explosion and subsequent ejecta-CSM interaction, becoming visible in the infrared (e.g. B. E. K. Sugerman, 2003; R. Kotak et al., 2009; O. D. Fox et al., 2010). Dust in nearby SNe is likely both pre-existing in the CSM and created in the ejecta. However, only a handful of SNe have the multi-epoch, multi-wavelength measurements crucial to disentangle the origin of the dust and therefore constrain the timeline of CCSNe dust formation.
The launch of JWST has ushered in a new era of SN dust studies. JWST’s sensitivity, resolution and wavelength coverage has allowed for observations that probe both hot and cool dust around the SNe in the years following collapse. In concert with multi-wavelength ground- and space-based observations, numerous studies have utilized JWST data to constrain the dust formation around some of the nearest SNe (R. G. Arendt et al., 2023; G. Hosseinzadeh et al., 2023b; O. C. Jones et al., 2023; M. Shahbandeh et al., 2023; P. Bouchet et al., 2024; S. Gomez et al., 2024; M. Matsuura et al., 2024; M. Shahbandeh et al., 2024a, b; S. Zsíros et al., 2024; G. C. Clayton et al., 2025; A. Sarangi et al., 2025; T. Szalai et al., 2025; S. Tinyanont et al., 2025). This growing dataset is vital to understanding the onset of dust production in CCSNe and its impact on the evolution of the early Universe.
In this paper, we present an analysis of newly acquired and archival multi-wavelength observations of SN 2017eaw, the first SN with multi-epoch late-time mid-infrared JWST observations. We review the observational data and reduction techniques in Section 2. In Section 3, we analyze the UV and infrared photometry, determine the extent of mid-infrared evolution, and model the dust SED. We discuss the possible origin of SN 2017eaw’s dust in Section 4 and conclude in Section 5.
2 Observations
SN 2017eaw was discovered on 2017 May 14 in NGC 6946, a nearby galaxy ( Mpc according to the latest TRBG distance222https://edd.ifa.hawaii.edu/get_cmd.php?pgc=65001, this distance is also used in M. Shahbandeh et al. (2023)) with a high SN rate (see Figure 1). SN 2017eaw has extensive multi-wavelength observations, both pre- and post-explosion (explosion on 57885.2 MJD S. D. Van Dyk et al., 2019), and has been the subject of numerous studies (e.g. D. Y. Tsvetkov et al., 2018; J. Rho et al., 2018; C. D. Kilpatrick & R. J. Foley, 2018; S. Tinyanont et al., 2019; S. D. Van Dyk et al., 2019; T. Szalai et al., 2019b; L. Rui et al., 2019; R. J. Buta & W. C. Keel, 2019; K. E. Weil et al., 2020; M. Shahbandeh et al., 2023; K. A. Bostroem et al., 2023).
There is significant evidence of ejecta-CSM interaction in SN 2017eaw. Pre-explosion Spitzer Space Telescope (hereafter Spitzer) images show the progenitor star was surrounded by a dusty shell at 4000 (C. D. Kilpatrick & R. J. Foley, 2018). Early time detections of SN 2017eaw in the X-ray, UV, and radio indicate moderate interaction between the SN shock and the CSM (T. Szalai et al., 2019b). Further, the optical light curve of SN 2017eaw exhibits a bump peaking around a week post-explosion that further suggests early time CSM interaction (T. Szalai et al., 2019b). Years after explosion there remain signs of ongoing ejecta-CSM interaction. Hubble Space Telescope (HST) near-UV imaging reveals the SN is still UV bright (as shown in Figure 1; S. D. Van Dyk et al., 2023) and late-time ( days post-explosion) optical spectra exhibit boxy line profiles indicating that the ejecta is continuing to collide with the surrounding material (K. E. Weil et al., 2020; M. Shahbandeh et al., 2023). Given the continued detection of CSM interaction, it is likely that some pre-existing dust surrounds SN 2017eaw.
In addition to the presence of dust in the CSM, there are several observational indicators that SN 2017eaw is producing dust in its ejecta. CO was detected in the near-infrared (NIR) spectra roughly one year post-explosion, demonstrating that the temperature of the ejecta has cooled enough for dust formation (J. Rho et al., 2018; S. Tinyanont et al., 2019). Further, nebular spectra of SN 2017eaw reveal blueshifted and asymmetric line profiles indicative of dust in the ejecta (J. Rho et al., 2018; K. E. Weil et al., 2020; M. Shahbandeh et al., 2023). The likely presence of both pre-existing and newly-formed dust in SN 2017eaw makes it an ideal test case for understanding when newly-formed dust begins to dominate the infrared dust spectral energy distribution (SED).
Post-explosion infrared observations of SN 2017eaw were executed by both Spitzer and JWST. Ground-based NIR and Spitzer observations (3.6 and 4.5 m) at 200 and 500 days post-explosion and JWST mid-infrared observations at days reveal a population of observed silicate dust that has increased in mass over the years (from to M⊙ S. Tinyanont et al., 2019; M. Shahbandeh et al., 2023). However, given the proximity of the Spitzer observations to explosion and the limited wavelength coverage, it is difficult to determine if the Spitzer and JWST dust populations are related.
We present the second epoch of mid-infrared JWST imaging of SN 2017eaw, taken one year after the first, and compare it directly to the previously published epoch in order to determine the nature and origin of the dust. To further understand the dust evolution of SN 2017eaw, we also collect and analyze additional new and archival data, including optical and near-ultraviolet HST observations and ground-based optical and infrared imaging and spectroscopy.
2.1 JWST/MIRI
Initial JWST Mid-Infrared Instrument (MIRI; P. Bouchet et al., 2015; G. H. Rieke et al., 2015; G. S. Wright et al., 2023) imaging of SN 2017eaw taken using the F560W, F1000W, F1130W, F1280W, F1500W, F1800W, F2100W, and F2550W filters was obtained in September 2022 (1957.7 days post-explosion) as part of the Cycle 1 General Observers (GO) 2666 Program (O. D. Fox et al., 2021). Photometry and analysis of this epoch was previously published in M. Shahbandeh et al. (2023).
Further JWST/MIRI observations of SN 2017eaw were also obtained on 26 September 2023 (2328.2 days post-explosion) as part of the Cycle 2 GO 3295 Program (D. J. Sand et al., 2023). These observations were taken with the complete set of MIRI filters (F560W, F770W, F1000W, F1130W, F1280W, F1500W, F1800W, F2100W, and F2550W), using the FULL array with a FASTR1 readout pattern, a 4-point dither pattern, and an exposure time of 111 seconds for all filters.
Given that one focus of this work is on the flux evolution between the Cycle 1 and 2 observations, we opt to reanalyze photometry of the Cycle 1 data so that the methodology remains consistent between epochs. In this work, both the Cycle 1 and Cycle 2 JWST observations of SN 2017eaw were processed with the JWST Calibration Pipeline version 1.15.1, with the Calibration Reference Data System version 11.17.25 (H. Bushouse et al., 2025).
We attempt aperture photometry on the Cycle 1 and 2 images using several different methods, outlined in Appendix A. However, M. Shahbandeh et al. (2023) report PSF photometry for SN 2017eaw, and we find that our aperture photometry methods result in flux values for the Cycle 1 observations which are 10-40% higher than those reported in M. Shahbandeh et al. (2023). For consistency with the published photometry, we instead report PSF photometry done using space_phot333space_phot version 0.2.5 https://space-phot.readthedocs.io (J. Pierel, 2024; J. D. R. Pierel et al., 2024). We note that this choice does not significantly impact the conclusions of this work since they are based primarily on the difference between the JWST/MIRI epochs and not the absolute flux calibration. PSF photometry with space_phot is done on the stage 2 products for all filters except F2550W. This involves fitting the SN’s PSF in each of the 4 individual Level 2 CAL files using WebbPSF (M. D. Perrin et al., 2012, 2014, version 1.2.1) models. Given the low signal to noise detection of SN2017eaw in F2550W, we opt to do PSF photometry on the Level 3 stacked images for this filter. The space_phot routine for Level 3 photometry uses temporally and spatially dependent Level 2 PSF models from WebbPSF, and drizzles them together to create a Level 3 PSF model. While the MIRI PSF models have been significantly updated since the publication of M. Shahbandeh et al. (2023), we find our Cycle 1 PSF photometry is mostly consistent with the previously reported values (see Appendix A for more information). We report the flux values from space_phot for both the Cycle 1 and Cycle 2 observations in Table 1.
| Cycle 1 | Cycle 2 | |||||||
| Filter | MJD | Phase444From explosion on 57885.2 MJD (S. D. Van Dyk et al., 2019) | Flux | AB Mag | MJD | Phase | Flux | AB Mag |
| [days] | [ mJy] | [days] | [ mJy] | |||||
| F560W | 59842.885 | 1957.7 | 60213.391 | 2328.2 | ||||
| F770W | – | – | – | – | 60213.394 | 2328.2 | ||
| F1000W | 59842.896 | 1957.7 | 60213.405 | 2328.2 | ||||
| F1130W | 59842.906 | 1957.7 | 60213.409 | 2328.2 | ||||
| F1280W | 59842.913 | 1957.7 | 60213.417 | 2328.2 | ||||
| F1500W | 59842.919 | 1957.7 | 60213.422 | 2328.2 | ||||
| F1800W | 59842.927 | 1957.7 | 60213.428 | 2328.2 | ||||
| F2100W | 59842.935 | 1957.7 | 60213.436 | 2328.2 | ||||
| F2550W | 59842.941 | 1957.7 | 60213.439 | 2328.2 | ||||
2.2 HST Optical and UV
Several HST observations of SN 2017eaw have been taken since explosion (see Table 2). S. D. Van Dyk et al. (2023) reported HST photometry of SN 2017eaw from late 2020 and early 2022 as part of their study of the progenitor. Since then observations in ACS/WFC F555W and F814W (PI: C. Kilpatrick, ID: 17070) and WFC3/UVIS F275W and F555W (PI: W. Jacobson-Galan, ID: 17506) in late 2022 and late 2023 respectively, have been completed.
We use DOLPHOT (A. E. Dolphin, 2000; A. Dolphin, 2016) to obtain PSF photometry of SN 2017eaw in all HST images. We use the calibrated and charge-transfer-efficiency (CTE) corrected flc and the corresponding drizzled drc frames from the Mikulski Archive for Space Telescopes (MAST) as inputs for DOLPHOT. Each epoch and filter combination was run through DOLPHOT separately and the flc frames were aligned to the associated drc image. We use the same DOLPHOT parameter settings as were used for the HST PHAT survey (J. J. Dalcanton et al., 2012; B. F. Williams et al., 2014). DOLPHOT detected a “good” star (“object type”=1) at the location of SN 2017eaw in all filters and epochs. Where available, we find our photometry is completely consistent with published values in S. D. Van Dyk et al. (2023). We present the PSF photometry of the detected source in Table 2 for SN 2017eaw. All UV, optical, and NIR magnitudes are reported in Vega magnitudes.
| Filter | MJD | Phase | Vega Mag | Tele/Inst |
| [days] | ||||
| F336W∗ | 59156.290 | 1271.1 | HST/WFC3 | |
| F275W∗ | 59156.389 | 1271.2 | HST/WFC3 | |
| F555W∗ | 59164.831 | 1279.6 | HST/WFC3 | |
| F814W∗ | 59164.813 | 1279.6 | HST/WFC3 | |
| F555W∗ | 59622.265 | 1737.1 | HST/WFC3 | |
| F814W∗ | 59622.259 | 1737.1 | HST/WFC3 | |
| F555W | 59924.463 | 2039.3 | HST/ACS | |
| F814W | 59924.457 | 2039.3 | HST/ACS | |
| F275W | 60299.132 | 2413.9 | HST/WFC3 | |
| F555W | 60299.141 | 2413.9 | HST/WFC3 | |
| K | 60040.505 | 2155.3 | MMT/MMIRS | |
| J | 60044.427 | 2159.2 | MMT/MMIRS | |
| H | 60044.488 | 2159.3 | MMT/MMIRS | |
| J | 60282.062 | 2396.9 | MMT/MMIRS | |
| K | 60282.092 | 2396.9 | MMT/MMIRS | |
| r† | 60203.144 | 2317.9 | MMT/Binospec | |
| i† | 60205.257 | 2320.1 | MMT/Binospec |
∗ Previously published in S. D. Van Dyk et al. (2023)
† AB magnitudes are r: and i:
2.3 MMT Optical and NIR Imaging
Additionally, we report optical and NIR ground-based photometry of SN 2017eaw. We present r and i band imaging of SN 2017eaw from 16 and 18 September 2023 (60203.144 and 60205.257 MJD) respectively, taken with the Binospec instrument on the MMT (D. Fabricant et al., 2019), and NIR JHK photometric observations of SN 2017eaw taken with the MMT and Magellan Infrared Spectrograph (MMIRS) on the MMT (B. McLeod et al., 2012) in Spring 2023 and December 2023.
For Binospec observations, we utilize a standard dither pattern. These images are then reduced using a custom python Binospec imaging reduction pipeline555Initially written by K. Paterson and available on GitHub: https://github.com/CIERA-Transients/POTPyRI, which does standard flat-fielding, sky background estimation, astrometric alignments, and stacking of the final individual exposures.
For MMIRS, each observation consisted of a dithered sequence which alternates between the target field and a off-galaxy field to allow for better sky subtraction given the IR-brightness of NGC6946. The resulting , , and band observations were reduced using a custom pipeline666Adapted from the MMIRS imaging pipeline available on github: https://github.com/CIERA-Transients/POTPyRI which does standard dark-current correction, flat-fielding, sky background estimation and subtraction, astrometric alignments, and stacking of the final individual exposures.
The total field-of-view (FOV) of MMIRS (6’.96’.9) and Binospec (two 8’15’ FOVs with 3’ gap) is large enough to calibrate photometric zeropoints using isolated stars with cataloged Two Micron All Sky Survey (2MASS; M. F. Skrutskie et al., 2006) and Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; K. C. Chambers et al., 2016), respectively. We derive an effective (e)PSF model for each image by fitting bright, isolated stars with the EPSFBuilder tool from the photutils package in Astropy. For all filters where the supernova is detected, we then perform PSF-fitting at the location of the target as well as a set of 20 or more stars spread throughout the image. A low-order, two-dimensional polynomial is included in the fit to account for any spatially varying background and avoid over-fitting of the stars. To estimate the statistical uncertainty of each flux measurement, we first set the statistical uncertainty per pixel using the RMS error of the fit residuals scaled by a factor of the square root of the reduced (usually 1), then multiply by the number of ‘noise pixels’ of the ePSF777A derivation of this quantity by F. Masci can be found here: http://web.ipac.caltech.edu/staff/fmasci/home/mystats/noisepix_specs.pdf. We use the set of 2MASS or Pan-STARRS calibration stars to derive aperture corrections (0.1 mag in all filters) to scale PSF-fitting magnitudes to the images’ photometric zeropoints. We adopt the statistical flux uncertainty summed in quadrature with the RMS error of the stars used in the zeropoint and ePSF aperture correction as the total uncertainty in our reported magnitudes. Despite the large FOV of both MMIRS and Binospec, the limited number of isolated 2MASS and Pan-STARRS stars means that the zeropoint RMS dominates the reported error.
SN 2017eaw was not detected in the MMIRS H-band and K-band images taken on 60044.427 and 60282.092 MJD (2159 and 2397 days post-explosion), respectively. For these observations, we instead report a 5 limiting magnitude, based off randomly placed background apertures near the location of SN 2017eaw. Optical (converted to Vega magnitudes) and NIR MMT photometry of SN 2017eaw are reported in Table 2.
2.4 Keck LRIS Spectroscopy
To complement the Cycle 2 JWST MIRI observations, we obtained an optical spectrum of SN 2017eaw on 2024 Aug 31 (60553.25 MJD, 2668 days post-explosion) using the Low Resolution Imaging Spectrometer (LRIS; J. B. Oke et al., 1995) on Keck I. The spectrum was taken with a 1.5” slit width with the 600/4000 grism and the 400/8500 grating at a central wavelength of 7700 Å and a total exposure time of 7200 seconds.
3 Analysis
We present the full time series spectral energy distribution (SED) of SN 2017eaw from 2020 to late 2023 in Figure 2. The flux has generally decreased in each consecutive epoch. The notable exception to this trend are the and bands, where the filter bandpasses include broad hydrogen, calcium, and oxygen lines which have been observed in nebular spectra of SN 2017eaw (see Section 3.7 and M. Shahbandeh et al., 2023). Strikingly, the mid-infrared fluxes blue-ward of 18 m have declined between Cycle 1 and Cycle 2 but are consistent between Cycle 1 and 2 for wavelengths m. As we discuss in Section 3.1 below, it is difficult to determine the statistical significance of the SED evolution redward of 15 m.
3.1 Comparison with Stars in the Field
To ensure that the observed decrease in luminosity is the result of a true decrease in flux and not the result of changes in the different observing parameters (i.e. exposure time, camera orientation, etc.) used in the Cycle 1 and Cycle 2 observations, we perform PSF photometry on several stars in the field using methods similar to those used for the SN 2017eaw photometry reported in Table 1.
We first identify bright IR objects with minimal variability in the field by using the 2MASS catalog (M. F. Skrutskie et al., 2006). These objects are then confirmed to be point sources in all of the JWST filters. Since star clusters may appear as point sources in MIRI filters, we also include a cut to remove any objects in crowded regions by using an HST F814W image of NGC6946 as a reference. This procedure results in 9 comparison stars for the SN 2017eaw field. Given the low signal to noise detections of these reference stars, particularly in the redder bands, we opt to do photometry on the stage 3 products for all filters rather than the stage 2 products as was done for SN 2017eaw. We note that stage 2 and 3 photometry of SN 2017eaw produce flux change measurements that are consistent within the uncertainties.
The percent difference in flux between the Cycle 1 () and Cycle 2 () comparison star observations is calculated as . To determine percent change in the total comparison sample we calculate the average change across the sample and adopt the standard deviation of the flux values as the error in this measurement.
As shown in Figure 3, we find that the comparison stars are consistent with no change in flux across the two epochs in the F560W, F770W, F1000W, F1130W, F1280W, and F1500W filters. The change in flux is consistent with zero redward of 15 m as well, with the exception of the F2550W filter. Filters where fewer reference stars are detected have larger errors due to small number statistics. For example, one of the reference stars has a high variance in F560W resulting in large error bars on the average flux change for this filter. Due to the high sky flux in the redder bands (18 m) there are significant errors in flux measurements for individual stars, large scatter between stars, and smaller sample sizes as few of our reference stars are detectable in the reddest bands. Therefore we can not make a high significance measurement of the extent of flux change in SN 2017eaw redward of 15 m.
3.2 Decrease in UV Flux
HST F275W and F336W observations from 2020 indicated that SN 2017eaw was UV bright. S. D. Van Dyk et al. (2023) suggested the elevated UV flux was the result of CSM-ejecta interaction but could not rule out the possibility of an underlying UV source like an O-star or small stellar cluster. As shown in Figure 2, the F275W observations taken in late 2023 reveal that, while the UV is still elevated, it has declined significantly. The F275W flux in 2023 is 16% of the F275W flux in 2020. Such a significant UV evolution is unlikely to be caused by an underlying star cluster or main sequence star.
It is notable that the F275W filter is particularly sensitive to CSM-ejecta interaction given it includes the Mg II 2796, 2803 doublet (L. Dessart et al., 2023). Boxy H emission lines, a signature of late time CSM interaction, have been observed in the optical spectra of SN 2017eaw since 900 days post-explosion (K. E. Weil et al., 2020), well before the 2020 F275W observations, and boxy H is still present in more recent spectra (see Section 3.7). Given the evolution and additional observational signatures, the UV evolution is very likely tracing the CSM-ejecta interaction.
3.3 Progenitor Disappearance
Given the proximity of the host galaxy, numerous images of the progenitor of SN 2017eaw were taken by both ground and space-based observatories. HST and Spitzer images identified the progenitor as a 12-15 M⊙ dusty red supergiant (C. D. Kilpatrick & R. J. Foley, 2018; S. D. Van Dyk et al., 2019; L. Rui et al., 2019). Since the supernova explosion, several epochs of HST photometry have been acquired. An analysis of the post-explosion imaging up to February 2022, suggested that the F814W flux had faded below the progenitor level (S. D. Van Dyk et al., 2023).
A further epoch of F814W imaging was obtained in December 2022 and we find it is similarly below the progenitor flux. Additionally, we compare the NIR progenitor flux to the MMT/MMIRS photometry from December 2023. As shown in Figure 4, we find that the and band detections are 2 magnitudes fainter than the progenitor. We are therefore able to confirm the progenitor identification and verify that it has significantly faded from pre-supernova observations.
3.4 Dust Modeling Methods
We assume the mid-infrared flux observed by JWST/MIRI in Cycle 1 and Cycle 2 is due to thermal emission from dust grains in or near the SN ejecta. To model the cool and hot dust components we employ several analytical dust models using a procedure adapted from G. Hosseinzadeh et al. (2023b).
Given the prominent shape of the silicate feature at m, the observed dust is unlikely to be optically thick. Therefore we first fit the dust using an optically thin model as was done in M. Shahbandeh et al. (2023) and S. Zsíros et al. (2024). We also model a dusty sphere (similar to M. Shahbandeh et al., 2023) and a dusty shell motivated by work presented in E. Dwek & R. G. Arendt (2024); for both of these dust models we allow the optical depth to vary. All of the models presented here assume there are two temperature components, as motivated by the SED shape, within the same geometry. Details of the luminosity equations for these three models can be found in Appendix B.
We fit all three models (Equations B3, B.2, and B12) to a filter-integrated model of the observed SED using the MCMC routine implemented in the Light Curve Fitting package (G. Hosseinzadeh et al., 2023a). Additionally, we fit an intrinsic scatter term, , which inflates the error bars on each photometric data point by a factor of , to account for underestimated photometric uncertainties, and to account for uncertainties in the model (e.g. infrared line emission). We run 20 walkers for 2000 steps to reach convergence and then 1000 more steps to properly sample the posterior. All of the optical filters blueward of 0.8 m include flux from broad emission lines (see Figure 2), so we exclude all of these filters when fitting the SED. We include only the NIR detections and the JWST/MIRI observations in our fit of the Cycle 2 (2023) SED. For Cycle 1, only the JWST/MIRI observations are considered.
As shown in Figure 2, the 5-25 m SED exhibits two distinct peaks with a trough at 13 m. This double humped shape is characteristic of optically thin silicate dust. Therefore we assume that the cool component is silicate dust, with g cm-3 and m as given by A. Laor & B. T. Draine (1993). The composition of the hot component is not as clearly identified. Recent studies report both amorphous carbon and silicate hot dust components around CCSNe (G. Hosseinzadeh et al., 2023b; M. Shahbandeh et al., 2023; S. Zsíros et al., 2024), so we try models with both silicate and amorphous carbon dust (m and from L. Colangeli et al., 1995) for the optically thin dust case. For the dusty sphere and dusty shell cases, we assume both components are silicate dust.
We attempt to fit a model of hot amorphous carbon dust and cool silicate optically thin dust, as was done for SN 1980K (S. Zsíros et al., 2024), to the NIR and JWST/MIRI photometry from Cycle 2. As shown in Figure 5, we find that the carbon+silicate model can not reproduce the photometry 15 m. We note that this excess could potentially be fit with an additional cooler component, as might be expected of dust formed in the ejecta or pre-existing dust at larger radii. However, given that the SED evolution redward of 15µm is not well constrained, we avoid implementing a third dust component. The hot carbon component requires a temperature of T K to reproduce the observed NIR peak, this is significantly above the condensation temperature of amorphous carbon dust (T K; K. Lodders & B. Fegley, 1997). We therefore deem it unlikely that the hot dust component in SN 2017eaw is primarily carbonaceous dust. However, we cannot rule out the possibility that there is both carbon and silicate dust in the ejecta, though the shape and temperature of the SED would suggest a high ratio of silicate to carbon dust if carbon dust is present.
Given the poorer quality fit of the carbon+silicate optically thin model, we focus primarily on models where both dust components are silicate. In Table 3, we list the model parameters, their priors, and their best-fit values (median and 1 equal-tailed credibility interval), for the optically thin, dusty sphere, and dusty shell models of the full considered SED for Cycle 1 and Cycle 2. The best-fit models for all three iterations, and the separate dust components, are shown in Figure 6. Since the Cycle 1 SED does not include F770W or NIR observations, we also fit the Cycle 2 SED excluding these filters. For completeness, we also present the results of the Cycle 2 fits excluding the NIR and F770W observations in Table 3 and discuss them further in Appendix C.
| Cycle 1 (1960 days) | Cycle 2 (2330 days) | ||||
| Parameter | Priors | All Filtersa | All Filtersb | No NIR & F770Wa | |
| Optically Thin | Thot | Uniform(1.0, 2.0) | |||
| Tcool | Uniform(0.05, 0.3) | ||||
| Mhot | Log-Uniform(1e-15, 1.) | ||||
| Mcool | Log-Uniform(1e-15, 1.) | ||||
| Intrinsic scatter | Gaussian(0., 20.) | ||||
| Dusty Sphere | Thot | Uniform(0.5, 2.0) | |||
| Tcool | Uniform(0.05, 0.3) | ||||
| Mhot | Log-Uniform(1e-15, 0.01) | ||||
| Mcool | Log-Uniform(1e-15, 0.01) | ||||
| Router | Log-Uniform(1, 2000.) | ||||
| Intrinsic scatter | Gaussian(0., 20.) | ||||
| Dusty Shell | Thot | Uniform(1.0, 2.0) | |||
| Tcool | Uniform(0.05, 0.3) | ||||
| Mhot | Log-Uniform(1e-10, 1.) | ||||
| Mcool | Log-Uniform(1e-8, 1.) | ||||
| Rinner | Log-Uniform(0, 2000.) | ||||
| Router | Log-Uniform(1, 5000.) | ||||
| Intrinsic scatter | Gaussian(0., 20.) | ||||
a includes same filters as Cycle 1 – full JWST/MIRI filter suite excluding F770W
b includes and in addition to full JWST/MIRI filter suite
3.5 Dust Modeling Results
Two silicate dust components are able to reproduce the complete SED for all three model geometries, as shown in Figure 6 (top: optically thin, middle: dusty sphere, bottom: dusty shell). While the F770W observation somewhat reduces the spread in the posterior distribution and constrains the hot dust component, the NIR photometry is the most significant factor in constraining the dust models given that the hot dust SED peaks near the effective wavelength of the filter. The particular dust model is not a significant factor in constraining the dust properties, all the models agree on the masses ( for the cool and for the hot dust) and temperatures ( K for the cool and K for the hot) of the components for either epoch. The inner radius in the dusty shell model is consistent with zero (i.e. a dusty sphere) for both Cycle 1 and 2 observations. Given this, we report the upper limit on Rinner in Table 3.
When we compare the Cycle 1 to the Cycle 2 full filter set models, we find that all of the best-fit values are consistent within error with the exception of the differences in Mhot (mass of hot dust component) for all three model geometries and Router for the dusty sphere model and dusty shell models. Given that the shape of the SED does not significantly change between Cycle 1 and Cycle 2, it is not unexpected that the temperatures of the dust components are roughly the same between epochs. If the dust is in the SN ejecta or pre-existing and actively interacting with the ejecta, we would expect some increase in the dust radius as the SN ejecta expands over the year between observations. Assuming an ejecta velocity of 7000 km/s (an upper estimate from the most recent spectrum, see Figure 8) in the 370 days between the Cycle 1 and Cycle 2 observations the ejecta radius should expand which, when accounting for errors, is consistent with the evolution between Router in Cycle 1 and Cycle 2 for the dusty sphere model and dusty shell models.
The difference in Mhot between epochs is likely due to the lack of constraints on the hot dust component in Cycle 1. Excluding F770W and the NIR photometry from the Cycle 2 fits produces Mhot values consistent with those observed in Cycle 1. This highlights the need for NIR photometry for constraining dust masses, particularly around SNe younger and hotter than SN 2017eaw. We note that there is no clear indication, in any of the three models, that the mass of the hot or cool dust components increased in the year between observations.
The Cycle 1 JWST observations were previously modeled for the optically thin and dusty sphere case in M. Shahbandeh et al. (2023). M. Shahbandeh et al. (2023) report only a total dust mass rather than separate mass components as we do here. However, we are able to reproduce all the dust properties for the dusty sphere of silicate dust case reported in M. Shahbandeh et al. (2023) by assuming since . For the optically thin case, we are unable to replicate the temperature of the hotter dust component as M. Shahbandeh et al. (2023), though we reproduce the mass of the dust (again assuming ). However, we find that our Cycle 1 value is consistent within error with our Cycle 2 value, which is well constrained by the NIR data.
3.6 Dust Geometry
One of the primary ways to distinguish between newly-formed and pre-existing dust is to compare the geometry of the dust shell to that of the ejecta. The majority of pre-existing dust will not survive the interaction with the forward shock and ejecta and therefore cannot be located within the ejecta. Therefore, if the dust shell is at a radius sufficiently interior to the outermost ejecta, the dust must be newly-formed. Similarly, if the dust is outside the outermost radius of the ejecta then it must be pre-existing.
First, to determine the robustness of any modeled dust radii measurements, we apply the same check as S. Zsíros et al. (2024) and calculate the optical depth as follows (L. B. Lucy et al., 1989):
where cm2 g-1 estimated for 0.1 m silicate dust (from grain properties in B. T. Draine & A. Li, 2007; A. Sarangi, 2022). In the case of optically thin dust, the minimum outer radius of the dust is set by the blackbody radius. The dusty sphere case produces the minimum blackbody radius which is and for Cycle 1 and Cycle 2 respectively. If we assume this radius and a total dust mass of (based on the dust mass from the optically thin dust model), Equation 1 gives . Therefore, the observable dust around SN 2017eaw may be optically thin.
We find the quality of the optically thin model fit to be comparable to that of the dusty sphere and dusty shell model fits for both epochs, further indicating that the dust may be optically thin. We also find that the dusty shell model converges in the case where the outer radius is set to be inside the ejecta radius (, i.e. the radius at 2300 days assuming a velocity of 7000 km/s) and also does so when the inner radius is set to be greater than the ejecta radius. Both of these scenarios result in similar dust masses and temperatures for both dust components. We, therefore, assume that the dust is optically thin enough that the radius of the dust cannot be well constrained. Nevertheless, we compare the measured dust radii in the dusty sphere and shell models to the ejecta for completeness.
The radius at which dust resides in the ejecta is often quoted as a velocity coordinate with respect to the ejecta in order to account for the fact that the ejecta, and anything within it, is expanding over time. The outer edge of the ejecta is at a velocity of km/s, given that the line profiles in the 1811 day and 2888 day nebular spectra indicate the outermost ejecta is interacting with CSM at this radius (see Section 3.7 and M. Shahbandeh et al., 2023). To determine the velocity coordinate of the dust, we assume the simplest case , where is the time since explosion, is the radius of the dust, and is the velocity. If we use the radius derived from the dusty sphere model (since ), we find that the velocity coordinate of the dust emission is km/s and km/s in Cycle 1 ( days) and Cycle 2 ( days), respectively.
The radius within which dust formation can begin is the subject of debate. Recent work by A. Sarangi (2022) suggests dust formation is confined to the 2500 km/s velocity coordinate, at least for the first 3000 days. Other studies have suggested this velocity coordinate may be as high as 5000 km/s (J. K. Truelove & C. F. McKee, 1999; K. Maguire et al., 2012). If we assume the radius from the dusty sphere model, the Cycle 1 dust component is near the region of the ejecta where dust formation may occur. The Cycle 2 dust component is consistent with dust both inside and outside the ejecta. For the dusty shell model, the velocity coordinate for Cycle 1 and 2 are and km/s respectively, both of which are consistent with dust located outside of the ejecta. Regardless of model, the radii from the dust modeling could account for both pre-existing and newly-formed dust geometries.
3.7 Spectral Evolution
As shown in Figure 7, the nebular spectrum of SN 2017eaw continues to exhibit a prominent broad boxy H profile, denoting continued CSM-ejecta interaction even at 2668 days after explosion. A boxy H profile was first observed in SN 2017eaw 900 days post explosion (K. E. Weil et al., 2020). A comparison of the 2668 day and 900 day H profiles reveals that the center of the 2668 day profile is perhaps somewhat more blueshifted, km/s compared to km/s, with a slightly steeper slope at the top of the line. This shape indicates dust attenuation, since the light from the receding ejecta, i.e. the red side of the line profile, is absorbed by the dust along the line of sight. However, it is difficult to robustly determine the impact of the dust attenuation given the signal to noise of the spectrum. The velocity of the ejecta, as measured at the location of full width half maximum, has decreased, 8000 km/s at 900 days compared to 7000 km/s at 2668 days, but some slowing is expected given the extent of the continued CSM interaction (L. Dessart, 2024). Ultimately, the H at 2668 days is remarkably similar to that at 900 days and minimal evolution seems to have occurred in almost 2000 days.
The three most prominent lines in the 2668 day Keck spectrum are H, [O II] , and [O III] . We compare the profiles of these lines in Figure 8. We treat the [O II] doublet as a line centered at 7324.5Å. A [Ca II] doublet can be present in the [O II] line complex but the calcium doublet is clearly subdominant to the [O II] lines at this epoch. Despite the existence of the [O III] doublet, the shape of the line profile in the 2668 day spectrum suggest the majority of the light is from the stronger of the lines, we therefore attribute the entire profile to [O III] . The [O III] line complex also contains H, which is visible on the blue shoulder. Again, [O III] is the stronger of the two lines. Therefore we treat the [O II] and [O III] lines as primarily oxygen in our analysis.
The edges of the oxygen profiles line up remarkably well with the edges of the H profile. The oxygen lines are notably attenuated on the red-side of the profile. This effect is less pronounced in the hydrogen line but still seems to be present. This might suggest there is a reservoir of dust inside the ejecta which is absorbing light from the far side of the supernova. If this attenuation is due to newly formed dust in the ejecta, rather than some asymmetry in the explosion and/or CSM, the strength of the attenuation and the low dust mass revealed in the JWST/MIRI images suggests that the majority of this interior dust is likely too optically thick to observe in the mid-infrared.
Recent work by L. Dessart et al. (2025) presents models of a SN II at 1000 days post-explosion that is CSM-interacting and also contains a small mass of dust in the cold dense shell. They find there is no significant dust attenuation effect for dust interior to the ejecta due to the small angle of the inner dust relative to the emitting region of the outer ejecta. Further, in contrast to non-interacting SNe, in the L. Dessart et al. (2025) CSM-interacting SN model interior dust has no impact on the line strengths, or the hydrogen-oxygen line ratios, due to the fact that 99.7% of the model emission is from the outer ejecta. Indeed, K. E. Weil et al. (2020) note no significant signs of dust attenuation in SN 2017eaw at 900 days post-explosion. At 2668 days post-explosion, the velocity of the ejecta (7000 km/s) is similar to that of the 1000 day model (8000 km/s) in L. Dessart et al. (2025) and it is possible that the observed line attenuation is not due to interior dust. Although this physical picture may still be valid, we caution against direct comparisons given that both the inner and outer regions have evolved for an additional 1700 days. The velocity of the ejecta at 2668 days suggests that the outer regions of the ejecta have somewhat slowed due to CSM interaction and the conclusions from 1000 days may no longer be valid. L. Dessart et al. (2025) also investigate the effect of , , and of dust in the ejecta at a velocity coordinate of 8000 km/s. These models produce H profiles very similar to the one at 2668 days. The model profiles include a dip in the center (near rest wavelengths) due to the increased optical depth of the limbs of the shell. We refrain from definitively linking a similar feature in the observed H profile with dust given the low SNR of the spectrum. Further modeling of the effects of dust in CSM-interacting SNe II at late times (1000 days post-explosion) is needed to understand the evolution of H and the strongest oxygen lines.
4 Dust Origin Scenarios
Despite the general decrease in mid-infrared flux from Cycle 1 to Cycle 2, the SED evolution of SN 2017eaw points to dust which has not significantly changed in the year between JWST/MIRI observations. Given the optically thin nature of the dust and the constant dust temperatures, this drop in flux could be due to a decline in dust mass. These observations indicate a possible trend to watch for in future observations and set a robust baseline against which future measurements can be compared. Even if the dust mass is not decreasing, it has not increased as might be expected if dust formation was actively occurring. Further, the lack of temperature evolution indicates that the mass of any newly-formed dust that cooled between the JWST observations is small compared to the total dust mass observable in the mid-infrared.
As was noted in M. Shahbandeh et al. (2023), even the cooler component of dust is too warm to be heated only by the ejecta of SN 2017eaw. The observed dust temperatures require an external heating mechanism to be present. Given that neither of the components have cooled or heated markedly since the initial observations reported in M. Shahbandeh et al. (2023), the external heating mechanism must be maintained over the course of the year between observations. The possible heating mechanisms depend on the location of the dust and can therefore be used to probe the dust origin. We explore three different scenarios where the mid-infrared dust is primarily 1) pre-existing in the CSM and collisionally heated; 2) pre-existing and radiatively heated, and 3) newly-formed in the ejecta and radiatively heated.
4.1 Collisionally Heated Pre-existing Dust
The mid-infrared SED of SN 2017eaw is somewhat similar in shape to that of SNe 1987A and 1980K at significantly later phases, see Figure 9. In the case of both SNe 1987A and 1980K, the 8-20 m emission is consistent with 150-180K silicate dust, remarkably similar to SN 2017eaw. Notably the mid-infrared flux in SN 2017eaw, which is primarily dominated by the cool dust component, is almost identical in shape to Spitzer IRS spectra of SN 1987A between 6000 and 8000 days post-explosion. This component in SN 1987A has been linked to the collisional heating of the equatorial ring (E. Dwek et al., 2010; R. G. Arendt et al., 2016, 2020), and a similar scenario was suggested for SN 1980K (S. Zsíros et al., 2024). The mass of dust in the 160K component of SN 2017eaw is 1 order of magnitude larger than observed in SN 1987A and 1 order of magnitude smaller than observed in SN 1980K.
To determine if it is plausible for the cooler component of SN 2017eaw’s mid-infrared dust to be collisionally heated via interaction between the ejecta and CSM, we follow the method of O. D. Fox et al. (2010) for estimating the mass of dust processed by the forward shock (see also O. D. Fox et al., 2011; S. Tinyanont et al., 2016; S. Zsíros et al., 2022, 2024). Any pre-existing dust must reside outside the evaporation radius, inside which the peak luminosity of the SN will have destroyed any pre-existing dust grains. Assuming the temperature and peak bolometric luminosity measured by T. Szalai et al. (2019b), 14,000K and 1043 erg/s (rounded from L erg/s) respectively, the evaporation radius (Revap) is R⊙. Revap is significantly less than the ejecta radius at 2,000 days. In both JWST epochs, the SN ejecta has far surpassed the evaporation radius and could feasibly be interacting with CSM containing pre-existing dust.
In the case of collisional heating, the hot, post-shocked gas heats a shell of pre-existing dust. The total mass of this dust can be determined from the volume of the emitting shell using equations for grain sputtering and by assuming a dust-to-gas ratio of 0.01 (O. D. Fox et al., 2010). This gives
where is the shock velocity, is the time since explosion, and is the dust grain size. Similar to S. Zsíros et al. (2024), we use 5,000 km s-1 and 15,000 km s-1 and 0.005 and 0.1 m (we assume m in our dust modeling) as our lower and upper bounds, respectively. Assuming these values, the range of dust masses that could be collisionally heated is . The total dust masses observed in both Cycle 1 and 2 are in the middle of this range. We note that the velocity of the ejecta measured from the nebular spectra, 7000 km s-1, with a 0.1 m dust grain could result in a collisionally heated dust mass of . This result suggests that the majority of the dust observed in the mid-infrared could reasonably be collisionally heated pre-existing dust.
The correlation between the UV and mid-infrared flux evolution might also suggest the observed dust is collisionally heated. As shown in Figure 10, the flux in F275W decreased 27% per year between 2020 and 2023. Similarly, the MIRI flux decreased an average of 21% (%) per year across all filters between the Cycle 1 (2022) and Cycle 2 (2023) observations. The F275W filter notably includes Mg II 2796, 2803, one of the UV features most strongly affected by CSM interaction (L. Dessart et al., 2023). Therefore the evolution in F275W can be used as a proxy for the extent of CSM. The drop in F275W flux and corresponding drop in the mid-infrared suggests that both wavelength regimes are probing the same medium. This correlation between UV and mid-infrared flux is similar to the X-ray (which similarly probes CSM interaction) and mid-infrared evolution observed in SN 1987A at 6000-8000 days, during which the IR-to-X-ray flux ratio remains constant. As shown in Figure 9, the mid-infrared SED of SN 2017eaw is nearly identical to that of SN 1987A during this epoch where the 8-20 m dust component is believed to be collisionally heated dust in the equatorial ring (E. Dwek et al., 2010; R. G. Arendt et al., 2016).
Furthermore, the reduction in UV flux does not correlate with a change in the temperature of the dust. If the dust, whether pre-existing or newly formed, is radiatively heated by CSM interaction, the temperature of the dust is expected to decrease with the UV flux and therefore CSM interaction. There is no evidence of the majority of the mid-infrared dust cooling between epochs.
4.2 Radiatively Heated Pre-existing Dust
The theory that the cool dust component is collisionally heated assumes a linear decline in UV luminosity from 2020 to 2023, but there is no UV data between these epochs to track the decline. It is possible that the UV luminosity was constant enough in the year between the JWST/MIRI observations for the dust temperature to not substantially change over the course of the year. In this scenario, radiative heating could still account for the lack of temperature evolution in the dust. The decrease in mid-infrared flux could be attributed to a changing geometry of the dust shell illuminated by the CSM interaction. We therefore can not use the similar UV and mid-infrared decline rates to completely rule out the possibility of radiative heating.
A simple IR echo model, assuming the light from the SN explosion excites pre-existing dust, places the echo radius at , where is the duration of the light echo (M. F. Bode & A. Evans, 1980; E. Dwek, 1983). Assuming a lower limit of days in order for the echo to still be detectable in the Cycle 2 mid-infrared observations, this gives , significantly above the outer dust radii given by the non-optically thin models in Section 3.4. When the ejecta is between the evaporation and the light echo radii, as is the case for SN 2017eaw, the luminosity from CSM-ejecta interaction heats the pre-existing grains creating a CSM echo (O. D. Fox et al., 2010, 2011)888Sometimes referred to as a circumstellar shock echo. M. Shahbandeh et al. (2023) found that for SN 2017eaw the optical luminosity necessary to heat the grains to the temperature of the cool dust component in the Cycle 1 data exceeds the observed optical luminosity. Unsurprisingly, we find this to be the case for the Cycle 2 dust as well.
However, L. Dessart & D. J. Hillier (2022) suggests that CSM-ejecta interaction may primarily produce UV emission, especially in Ly and Mg II 2796, 2803. Assuming constant luminosity across the UV ( nm), we use the F275W observation to estimate in 2023. This may be an overestimation of the total UV luminosity given the Mg II lines fall into the F275W filter. Nevertheless, this value is similar to the modeled in L. Dessart & D. J. Hillier (2022), indicating that the UV luminosity could account for the temperature of the observed mid-infrared dust. More extensive wavelength coverage in the UV is required to place robust limits on the UV luminosity.
4.3 Newly-formed Dust
Pre-existing dust does not preclude the existence of newly-formed dust in the ejecta. The detection of CO in the ejecta between days post explosion indicates that the ejecta has long been cool enough to form dust (S. Tinyanont et al., 2019). Further, there are signs of blue-shifted line profiles in the spectra at roughly 2000 days (just before the JWST Cycle 1 observations) (M. Shahbandeh et al., 2023) and at 2668 days (as discussed in Section 3.7). The observed dust attenuation in the red side of the line profiles suggests there is dust in the ejecta of the SN.
The optical spectra of SN 2017eaw indicate that the SN is producing dust (although see L. Dessart et al., 2025). However, the newly-formed dust might be too optically thick to be observed in the mid-infrared even at days post explosion. In this case, only thermal emission from the outermost shell of the total mass of newly-formed dust will be observable. This outermost layer would only constitute an extremely small percentage of the total newly-formed dust mass and may not noticeably change the mid-infrared SED if a more massive amount of pre-existing dust is also present. Nevertheless, geometrical arguments (see Section 3.6) indicate that some of the observed mid-infrared dust could be located within the ejecta or in a cold dense shell between the forward and reverse shock. Given the optical depth and evolution of the cool dust component, it is unlikely to be primarily newly-formed though some component of this dust may be.
In contrast, there are minimal constraints on the evolution of the hot dust component due to the lack of NIR observations during Cycle 1, and there may have been dust growth and/or cooling between the Cycle 1 and 2 observations as would be expected of a newly formed dust component. There is no NIR spectra of SN 2017eaw at this late epoch so we are unable to quantify the extent to which the NIR emission is from the SN itself rather than dust. NIR spectra of SN 1987A at around 2000 days shows strong emission lines in band but none in band (A. Fassia et al., 2002). If this is also the case for SN 2017eaw, the hot dust component could be slightly cooler than modeled. Further, there is likely some flux in this component that is due to blackbody emission from the SN. Our inferred cool dust mass of (see Table 3) should be treated as a upper limit of the amount of dust found in the hot component. Given the small amount dust in the hot component (relative to the cold component), this uncertainty does not impact the conclusions of this work. Ultimately, we are unable to confirm the origin of the hot dust component from the existing observations. Further observations of SN 2017eaw should also include NIR observations to place constraints on the evolution of the hot component and provide insights into its origin.
4.4 Implications for Dust Formation in CCSNe
Significant work has been done to understand the timeline of dust formation in CCSNe. When SN 2017eaw’s mid-infrared dust mass is compared to literature values of CCSNe dust masses, it lies near the lower limit of the dust trend observed in M. Niculescu-Duvaz et al. (2022), as shown in Figure 11. The slight fluctuation in dust mass from to days post-explosion is similar to trends observed in several other SNe, though this behavior has never been observed in another SN days post-explosion. However, SN 2017eaw is the only SN other than SN 1987A with multiple epochs of mid-infrared observations between 1000-5000 days post-explosion.
If the mid-infrared dust emission in SN 2017eaw is primarily due to pre-existing dust, then its location on the dust formation timeline may be significantly different than shown in Figure 11. It is possible that many of the early time mid-infrared dust measurements of CCSNe are similarly contaminated with pre-existing dust. Late time dust mass measurements of SN 1987A were done in the far-IR and sub-mm and probed dust significantly colder than can be observed with JWST (M. Matsuura et al., 2011; R. Indebetouw et al., 2014; M. Matsuura et al., 2015). For SN 1980K, reported dust measurements were measured by modeling the dust attenuation on optical spectral lines. The mid-infrared SED of SN 1980K reveals 100 times less dust than indicated by the line profiles (S. Zsíros et al., 2024). This suggests that the majority of the dust in SN 1980K is also too cold to be observed by JWST/MIRI. The same might be true for SN 2017eaw but the signal to noise of the recent spectra makes modeling of the line profiles difficult.
However, SN 2017eaw is significantly younger than both SN 1980K and SN 1987A. Any newly-formed dust around SN 2017eaw should be more optically thick than observed in the two older supernovae. In the case of optically thick dust, radiative processes from CSM-ejecta interaction will only heat the outermost shell of newly-formed dust, which is likely to make up only a tiny amount of the total dust mass. Over the course of a year, the expansion of a shell of newly-formed dust may not be enough to visibly evolve the mid-infrared SED.
5 Summary & Conclusion
We present late time UV, optical, and near-infrared observations of SN 2017eaw to map its spectral energy evolution. The SN has declined in flux across almost all wavelengths. We find that the NIR flux has declined below the progenitor level, confirming the progenitor detection. SN 2017eaw is still detected in HST WFC3/UVIS F275W, and the optical spectrum at 2668 days exhibits broad boxy line profiles, particularly H, indicating that there is continued CSM-ejecta interaction even at 2500 days post-explosion.
SN 2017eaw is one of the first supernovae to have multi-epoch JWST MIRI imaging. These observations reveal that the mid-infrared flux has decreased in most filters in the year between the MIRI observations. SED modeling reveals a hot (1700 K) silicate dust component of and a cool (160 K) silicate dust component of consistent with being optically thin. Here we cite the dust modeling values for the dusty shell case since these values are between the optically thin and dusty sphere models. Interestingly, there is no indication that the dust is cooling or increasing in mass as might be expected for dust which is actively forming.
Furthermore, the decline in mid-infrared flux is similar to that observed in the UV, perhaps hinting that the dust observed in the mid-infrared is located in the same CSM whose interaction with the ejecta is producing UV flux. The evolution in the UV suggests a changing CSM density or geometry around SN 2017eaw. To understand how this continues to affect the dust budget and the dust heating mechanism, continued X-ray and UV observations are necessary.
The multi-wavelength evolution of SN 2017eaw suggest that, while there may be newly-formed dust in the ejecta or cold dense shell, a significant fraction of the cool dust observed in the mid-infrared is likely pre-existing. There is a need for further late time (1000 days post-explosion) multi-wavelength observations for the nearest supernovae, like SN 2017eaw, in order to map the extent and duration of CSM-interaction and its impact on dust evolution. Such observations, spanning the UV to the mid-infrared, will reveal insights into red supergiant mass loss in the final years before death and help to constrain the timeline of new dust production in SNe II.
6 Acknowledgments
We thank R.G. Arendt for providing the Spitzer IRS spectra of SN 1987A. Thank you to C. DeCoursey and J. Pierel for their help with JWST photometric reduction using space_phot. We also thank D. Perley for helpful advice in reducing the SN2017eaw Keck LRIS spectrum.
This work is based in part on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program GO3295 and GO2666.
This research is based in part on observations made with the NASA/ESA Hubble Space Telescope obtained from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-26555. These observations are associated with programs SNAP17070 and SNAP17506.
Observations reported here were obtained at the MMT Observatory, a joint facility of the University of Arizona and the Smithsonian Institution.
Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation.
The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain.
JEA is supported by the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America.
Time domain research by the University of Arizona team and D.J.S. is supported by National Science Foundation (NSF) grants 2108032, 2308181, 2407566, and 2432036 and the Heising-Simons Foundation under grant 2020-1864.
CDK gratefully acknowledges support from the NSF through AST-2432037, the HST Guest Observer Program through HST-SNAP-17070 and HST-GO-17706, and from JWST Archival Research through JWST-AR-6241 and JWST-AR-5441.
W.J.-G. is supported by NASA through Hubble Fellowship grant HSTHF2-51558.001-A awarded by the Space Telescope Science Institute, which is operated for NASA by the Association of Universities for Research in Astronomy, Inc., under contract NAS5-26555.
AAM, CL, and NR are supported by DoE award no. DE-SC0025599. MMT Observatory access for AM, CL, and NR was supported by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA).
KAB is supported by an LSST-DA Catalyst Fellowship; this publication was thus made possible through the support of Grant 62192 from the John Templeton Foundation to LSST-DA.
Appendix A Apperture vs. PSF Photometry
In order to measure the flux in the MIRI images, we attempted several different photometric techniques which we compared to the published Cycle 1 photometry in M. Shahbandeh et al. (2023). In this work, we report JWST MIRI PSF photometry of SN 2017eaw as this methodology resulted in values that are most consistent with previously published photometry. M. Shahbandeh et al. (2023) did PSF photometry on the stage 2 products, a method similar to the one we use for our photometry in Section 2.1. As shown in Figure 12 (purple diamonds and pink squares), the most significant offsets between our space_phot PSF photometry and the published photometry are at 18 and 21 µm, both of which are filters where the photometric calibration maps were significantly updated999This work uses version 0056, details can be found at https://jwst-crds.stsci.edu/browse/jwst_miri_photom_0056.rmap between the publication of M. Shahbandeh et al. (2023) and the completion of this work, therefore this offset is unsurprising.
Aperture photometry methods were unable to reproduce the flux values measured by PSF photometry methods. Our initial photometry of SN 2017eaw was done using an aperture photometry method similar to that used in G. Hosseinzadeh et al. (2023b), on the Cycle 1 and 2 Level 3 I2D images of SN 2017eaw. The science and background apertures for a selection of the images for SN 2017eaw is shown in Figure 13. We choose the science aperture to enclose 60% of the light from the source based on the JWST/MIRI aperture correction (version 0014, in flight pedigree of 2022-05-25 to 2024-06-02). Background subtraction is done using the average of two circular regions on either side of the aperture. We find that in the case of SN 2017eaw the diffraction spikes are minimal enough that using an annulus for background subtraction produces photometry which is consistent with that from the two circular aperture method. However, given the diffraction spikes and surroundings of the comparison stars in the field, we use two circular apertures for consistency. The location of the two background regions were chosen to avoid diffraction spikes while remaining close to the science aperture. The background apertures are chosen to be on the same region of the sky for all exposures of the target. However, as shown in Figure 12 (green circles), this methodology results in fluxes that are systematically higher than those published in M. Shahbandeh et al. (2023).
Given the discrepancy between the aperture photometry and the M. Shahbandeh et al. (2023) values, in order to cross check our methodology we utilize a separate aperture photometry code originally designed to do photometry on NIRCAM high-redshift galaxies described in R. Endsley et al. (2023), which was adapted to allow for aperture photometry on MIRI images. First, SEP (the python library for Source Extraction and Photometry; E. Bertin & S. Arnouts, 1996; K. Barbary, 2016) is run on the F2100W image to choose an elliptical aperture which encloses % of the flux and use this aperture size for all filters. We choose F2100W since it is the reddest filter with a high signal-to-noise point source at the location of SN 2017eaw, and therefore has the largest PSF. Second, SEP is run on all filters to mask out all objects in the field. This requires a background subtracted image. Given the complex background of the image, we opt to create 10”x10” stamps centered around the SN position and measure the spatially varying background using SEP. This background is then subtracted from the image stamp. Third, we randomly place 50 apertures on the background subtracted image and measured their fluxes. The standard deviation in these measurements is the error in our photometry and the median value is subtracted off the final photometry to account for higher order background fluctuations. Finally, we apply an aperture correction determined by dropping the elliptical aperture used on the WebbPSF models (M. D. Perrin et al., 2012, 2014) for each filter and calculating the amount of total flux enclosed within the aperture.
As shown in Figure 12 (blue stars), the random aperture background method also yields flux values higher than those reported in M. Shahbandeh et al. (2023). However, because of how photometric errors are accounted for in the random aperture background method, the error on these measurements are large. Therefore the photometry from this method is roughly consistent with both the previously discussed aperture photometry method and the M. Shahbandeh et al. (2023) PSF photometry.
Due to the discrepancy between the aperture and PSF photometry regardless of methodology, and the existence of published JWST/MIRI PSF photometry of SN 2017eaw, we opt to report only the PSF photometry discussed in Section 2.1 in this work. Importantly, we do find that, regardless of the photometric method, the total mid-infrared flux of SN 2017eaw has decreased from Cycle 1 to Cycle 2. We caution that the aperture and PSF photometry of JWST/MIRI data may not be consistent with each other and recommend using similar methods as reported in previous publications. This discrepancy may decrease over the course of the JWST mission as MIRI aperture corrections and PSF models continue to be updated.
Appendix B Dust Modeling Equations
In Section 3.4, we discuss the fitting methods for the optically thin, dusty sphere, and dusty shell models. Here we present the luminosity equations used for all three of these dust models.
B.1 Optically Thin Model
For the optically thin case, we model the input luminosity as two components of dust, with temperatures and and masses and . We note that in the optically thin case, Rdust is not well constrained. For each component, the input luminosity is set by:
where Bν(T) is the Planck function and is the frequency-dependent opacity of the dust component. We calculate from the absorption efficiency , particle density , and particle size :
In the optically thin case, the dust will not self attenuate so the total luminosity is just:
B.2 Dusty Sphere Model
In the dusty sphere case, we assume a sphere of dust with total mass and two temperature components ( and ) inside a radius with an optical depth . This model is geometrically similar to the dusty sphere model used in M. Shahbandeh et al. (2023). The luminosity of the dusty sphere is extinguished according to the escape probability from D. E. Osterbrock (1989, Appendix 2):
here the frequency dependent optical depth () to the center of a dust shell with bulk density is:
For the dusty sphere model is:
Thus the full SED for the dusty sphere is modeled by:
B.3 Dusty Shell Model
We also fit a dusty shell model, with total dust mass , inner radius and outer radius , to the SED. For the dusty shell, the frequency dependent optical depth (analogous to Equation B5) is:
The escape probability is similarly more complex, as it must take the inner cavity, where , into account. We use the escape probability expression worked out in E. Dwek & R. G. Arendt (2024):
where
and
where , , and and are the inner and outer radii of the shell, respectively.
Which yields the total dusty shell luminosity:
Appendix C Cycle 2 Dust models excluding F770W and NIR Observations
There was no NIR or F770W data taken coincident with the Cycle 1 observations of SN 2017eaw. The dust modeling presented here highlights the need for additional constraints of the hot component of the dust SED, especially for younger SNe like SN 2017eaw. Given the lack of constraints on the hot component at 1960 days, we also fit the 2330 day SED with the NIR and F770W photometry excluded so that the SEDs are directly comparable. These results are presented in the furthest right column of Table 3 and fits are displayed in Figure 14.
We find that the dust models without the NIR and F770W observations are consistent with the values determined for Cycle 1. The Cycle 2 models that exclude NIR and F770W photometry tend to favor lower temperatures for the hot components than those found for Cycle 1. Given that the hotter dust component is set by only F560W in these fits, we attribute the decrease in temperature to the decrease in F560W flux. Without NIR observations, it is impossible to track the temperature evolution of the hot dust component. This uncertainty highlights the need for NIR observations to complement further SNe dust studies. In the case of SN 2017eaw, the majority of the dust mass is in the cool component and the MIRI observations alone are sufficient to suggest that the dust mass of SN 2017eaw did not markedly increase over the course of a year. However, we caution that this may not be the situation for every supernovae, therefore NIR observations are crucial to understanding the dust budget of core-collapse supernovae.